OpenStudy (anonymous):

Which answer shows this equation in standard form? 8 – 4(x – y) = –6x + 6 A. 2x + 4y = 2 B. –10x + 4y = –2 C. 2x + 4y = –2 D. 2x – 4y = –2

OpenStudy (anonymous):

@Taylor<3sRin

OpenStudy (anonymous):

Explane how you got your answer :) thank you

OpenStudy (anonymous):

@Jackky

OpenStudy (jackky):

mija imma try my hardest

OpenStudy (jackky):

@AlexisPooBear

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

yes

OpenStudy (karon86):

I can do it. Would you like my help?

OpenStudy (jackky):

ay mija im so sorry

OpenStudy (jordanloveangel):

8 - 4(x - y) = 2x + 6 8 - 4x +4y = 2x + 6 2 = 6x -4y

OpenStudy (jhannybean):

Standard form of the equation of a line : \(\sf ax+by=c\)

OpenStudy (jordanloveangel):

6x - 4y = 2

OpenStudy (jhannybean):

So you want to move all your variables on one side, and keep all your constants on the other.

OpenStudy (jordanloveangel):

@Karon86 if u need help tag me

OpenStudy (jhannybean):

\[\sf 8 – 4(x – y) = –6x + 6\]Expand :\(-4(x-y)\) first. Combine all your x terms and your y terms on the left side, and subtract \(\sf -8\) from both sides. Then just simplify.

OpenStudy (anonymous):

how long i've been gone -.- lol